OpenStudy (anonymous):

Simplify: square root of 300 x to the sixteenth power

OpenStudy (alexandervonhumboldt2):

\[\sqrt{300}^{16}=\sqrt{100*3}^{16}=10\sqrt{3}^{16}=\]

OpenStudy (alexandervonhumboldt2):

woops forgot about x

OpenStudy (alexandervonhumboldt2):

\[10\sqrt{3x}^{16}=10*3x^{14}=47829690x^{14}\]

alones (alones):

i thought this equation cannot be simplified

OpenStudy (alexandervonhumboldt2):

i probably made mistake in math

OpenStudy (welshfella):

do you mean the sqrt(300x) or (sqrt300)x